You are asked to evaluate the following two projects for the Norton corporation. Use a discount rate of 10 percent. Use Appenc. Which project would you select based on the profitability index? O Project X O Project Y
Answer #1
  1. Profitability Index for Project X

Profitability Index = Present Value of Future Cash Flows ÷ Initial Investment

Initial Investment = $24,000

Present Value of Future Cash Flows

Year

Cash Flows In $

Present value factor @ 10%

Present Value

(I)

(II)

(III)

(II) * (III)

1

$12,000

0.9091

$10,909.20

2

$10,000

0.8264

$8,264.00

3

$11,000

0.7513

$8,264.30

4

$10,600

0.6830

$7,239.80

Present Value of the future Cash inflows

$34,677.30

Profitability Index = $34,677.30 ÷ $24,000

                                    =1.44

  1. Profitability Index for Project Y

Profitability Index = Present Value of Future Cash Flows ÷ Initial Investment

Initial Investment = $44,000

Present Value of Future Cash Flows

Year

Cash Flows In $

Present value factor @ 10%

Present Value

(I)

(II)

(III)

(II) * (III)

1

$22,000

0.9091

$20,000.20

2

$15,000

0.8264

$12,396.00

3

$16,000

0.7513

$12,020.80

4

$18,000

0.6830

$12,294.00

Present Value of the future Cash inflows

$56,711.00

Profitability Index = $56,711 ÷ $44,000

                                    =1.29

  1. Which Project to be selected
  • Project A

Decision Rule

IF

Decision

Profitability Index > 1

Accept the Project

Profitability Index < 1

Reject the project

In the given case both project is having Profitability Index > 1, which mean both project can be accepted. Since project A Profitability Index higher than project B, so we prefer project A.

Profitability Index is higher the better.

Note:

Calculation of Present value Factor

Discount Factor = 1/ (1+R) N

R = Discount Rate (i.e. = 10%)

N = No of years

E.g. for year 2 Discount Factor = 1/ (1.10)2

                                                                = 1/ (1.10) (1.10)

                                                = 0.8264